When Maria was born, her mother was 24 years old. At present, Maria's age is 20% of her mother's age. How old is Maria now?
Answer by stanbon(75887)

When Maria was born, her mother was 24 years old. At present, Maria's age is 20% of her mother's age. How old is Maria now?
---------
m = M + 24
M = 0.2m
------
Substitute for "M" and solve for "m".
m = 0.2m + 24
0.8m = 24
m = (10/8)24
m = 30 yrs (mother's age now)
-----
Solve for "M"::
m = M+24
30 = M + 24
M = 6 yrs (Maria's age now)
============================
Cheers,
Stan H.
==============